1.) If cos (40+x) = sin30, find x.
2.)If sectheta (1+sin theta ) (1-sin theta) = k, find k.
 

Grade:10

1 Answers

Ns
6 years ago
1 = ....x = 20 .....as Cos(40+x)= sin(90-(40+x))Therefore ........Sin(90-40-x) =sin 30=>50-x=30Therefore x = 20